# Model Lakes

### v1

@@ Acknowledgements.
- Mitchell et al. (2019) Margaret Mitchell, Simone Wu, Andrew Zaldivar, Parker Barnes, Lucy Vasserman, Ben Hutchinson, Elena Spitzer, Inioluwa Deborah Raji, and Timnit Gebru. 2019. Model Cards for Model Reporting. In Proceedings of the Conference on Fairness, Accountability, and Transparency (Atlanta, GA, USA) (FAT* ’19). Association for Computing Machinery, New York, NY, USA, 220–229. https://doi.org/10.1145/3287560.3287596
@@
- Mu et al. (2023) Xin Mu, Yu Wang, Yehong Zhang, Jiaqi Zhang, Hui Wang, Yang Xiang, and Yue Yu. 2023. Model Provenance via Model DNA. arXiv:2308.02121 [cs.LG]
@@
- OpenAI ([n.d.]a) OpenAI. [n.d.]a. Chat GPT. https://chat.openai.com/chat.
- OpenAI ([n.d.]b) OpenAI. [n.d.]b. Introducing GPTs. https://openai.com/blog/introducing-gpts.
@@
- Ramesh et al. (2022) Aditya Ramesh, Prafulla Dhariwal, Alex Nichol, Casey Chu, and Mark Chen. 2022. Hierarchical Text-Conditional Image Generation with CLIP Latents. arXiv:2204.06125 [cs.CV]
@@
- Ribeiro et al. (2016) Marco Tulio Ribeiro, Sameer Singh, and Carlos Guestrin. 2016. ” Why should i trust you?” Explaining the predictions of any classifier. In Proceedings of the 22nd ACM SIGKDD international conference on knowledge discovery and data mining. 1135–1144.
- Rombach et al. (2022) Robin Rombach, Andreas Blattmann, Dominik Lorenz, Patrick Esser, and Björn Ommer. 2022. High-resolution image synthesis with latent diffusion models. In Proceedings of the IEEE/CVF conference on computer vision and pattern recognition. 10684–10695.
@@
- Saharia et al. (2022) Chitwan Saharia, William Chan, Saurabh Saxena, Lala Li, Jay Whang, Emily Denton, Seyed Kamyar Seyed Ghasemipour, Burcu Karagol Ayan, S. Sara Mahdavi, Rapha Gontijo Lopes, Tim Salimans, Jonathan Ho, David J Fleet, and Mohammad Norouzi. 2022. Photorealistic Text-to-Image Diffusion Models with Deep Language Understanding. arXiv:2205.11487 [cs.CV]
@@
- Singh et al. (2024) Chandan Singh, Jeevana Priya Inala, Michel Galley, Rich Caruana, and Jianfeng Gao. 2024. Rethinking Interpretability in the Era of Large Language Models. arXiv:2402.01761 [cs.CL]
@@
- Touvron et al. (2023) Hugo Touvron, Louis Martin, Kevin Stone, Peter Albert, Amjad Almahairi, Yasmine Babaei, Nikolay Bashlykov, Soumya Batra, Prajjwal Bhargava, Shruti Bhosale, Dan Bikel, Lukas Blecher, Cristian Canton Ferrer, Moya Chen, Guillem Cucurull, David Esiobu, Jude Fernandes, Jeremy Fu, Wenyin Fu, Brian Fuller, Cynthia Gao, Vedanuj Goswami, Naman Goyal, Anthony Hartshorn, Saghar Hosseini, Rui Hou, Hakan Inan, Marcin Kardas, Viktor Kerkez, Madian Khabsa, Isabel Kloumann, Artem Korenev, Punit Singh Koura, Marie-Anne Lachaux, Thibaut Lavril, Jenya Lee, Diana Liskovich, Yinghai Lu, Yuning Mao, Xavier Martinet, Todor Mihaylov, Pushkar Mishra, Igor Molybog, Yixin Nie, Andrew Poulton, Jeremy Reizenstein, Rashi Rungta, Kalyan Saladi, Alan Schelten, Ruan Silva, Eric Michael Smith, Ranjan Subramanian, Xiaoqing Ellen Tan, Binh Tang, Ross Taylor, Adina Williams, Jian Xiang Kuan, Puxin Xu, Zheng Yan, Iliyan Zarov, Yuchen Zhang, Angela Fan, Melanie Kambadur, Sharan Narang, Aurelien Rodriguez, Robert Stojnic, Sergey Edunov, and Thomas Scialom. 2023. Llama 2: Open Foundation and Fine-Tuned Chat Models. arXiv:2307.09288 [cs.CL]
@@
- Vartak et al. (2018) Manasi Vartak, Joana M. F. da Trindade, Samuel Madden, and Matei Zaharia. 2018. MISTIQUE: A System to Store and Query Model Intermediates for Model Diagnosis. In Proceedings of the 2018 International Conference on Management of Data (Houston, TX, USA) (SIGMOD ’18). Association for Computing Machinery, New York, NY, USA, 1285–1300. https://doi.org/10.1145/3183713.3196934
- Vaswani et al. (2017) Ashish Vaswani, Noam Shazeer, Niki Parmar, Jakob Uszkoreit, Llion Jones, Aidan N. Gomez, Łukasz Kaiser, and Illia Polosukhin. 2017. Attention is all you need. In Proceedings of the 31st International Conference on Neural Information Processing Systems (Long Beach, California, USA) (NIPS’17). Curran Associates Inc., Red Hook, NY, USA, 6000–6010.
@@
- Zhao et al. (2021) Yan Zhao, Franck Ravat, Julien Aligon, Chantal Soule-dupuy, Gabriel Ferrettini, and Imen Megdiche. 2021. Analysis-oriented Metadata for Data Lakes. In Proceedings of the 25th International Database Engineering & Applications Symposium (Montreal, QC, Canada) (IDEAS ’21). Association for Computing Machinery, New York, NY, USA, 194–203. https://doi.org/10.1145/3472163.3472273
- Zheng et al. (2023) Lianmin Zheng, Wei-Lin Chiang, Ying Sheng, Siyuan Zhuang, Zhanghao Wu, Yonghao Zhuang, Zi Lin, Zhuohan Li, Dacheng Li, Eric. P Xing, Hao Zhang, Joseph E. Gonzalez, and Ion Stoica. 2023. Judging LLM-as-a-judge with MT-Bench and Chatbot Arena. arXiv:2306.05685 [cs.CL]
- Zhu et al. (2020) Chen Zhu, Ankit Singh Rawat, Manzil Zaheer, Srinadh Bhojanapalli, Daliang Li, Felix Yu, and Sanjiv Kumar. 2020. Modifying Memories in Transformer Models. arXiv:2012.00363 [cs.CL]
@@
- Zou et al. (2023) Andy Zou, Long Phan, Sarah Chen, James Campbell, Phillip Guo, Richard Ren, Alexander Pan, Xuwang Yin, Mantas Mazeika, Ann-Kathrin Dombrowski, et al. 2023. Representation engineering: A top-down approach to ai transparency. arXiv preprint arXiv:2310.01405 (2023).

### v2

@@ Acknowledgements.
- Mineault et al. (2024) Patrick Mineault, Niccolò Zanichelli, Joanne Zichen Peng, Anton Arkhipov, Eli Bingham, Julian Jara-Ettinger, Emily Mackevicius, Adam Marblestone, Marcelo Mattar, Andrew Payne, Sophia Sanborn, Karen Schroeder, Zenna Tavares, and Andreas Tolias. 2024. NeuroAI for AI Safety. arXiv:cs.AI/2411.18526 https://arxiv.org/abs/2411.18526
@@
- Mitchell et al. (2019) Margaret Mitchell, Simone Wu, Andrew Zaldivar, Parker Barnes, Lucy Vasserman, Ben Hutchinson, Elena Spitzer, Inioluwa Deborah Raji, and Timnit Gebru. 2019. Model Cards for Model Reporting. In Proceedings of the Conference on Fairness, Accountability, and Transparency (FAT* ’19). Association for Computing Machinery, New York, NY, USA, 220–229. https://doi.org/10.1145/3287560.3287596
@@
- Morris et al. (2023) John Morris, Volodymyr Kuleshov, Vitaly Shmatikov, and Alexander Rush. 2023. Text Embeddings Reveal (Almost) As Much As Text. In Proceedings of the 2023 Conference on Empirical Methods in Natural Language Processing, Houda Bouamor, Juan Pino, and Kalika Bali (Eds.). Association for Computational Linguistics, Singapore, 12448–12460. https://doi.org/10.18653/v1/2023.emnlp-main.765
- Morris et al. (2024) John Xavier Morris, Wenting Zhao, Justin T Chiu, Vitaly Shmatikov, and Alexander M Rush. 2024. Language Model Inversion. In The Twelfth International Conference on Learning Representations. https://openreview.net/forum?id=t9dWHpGkPj
- Mu et al. (2023) Xin Mu, Yu Wang, Yehong Zhang, Jiaqi Zhang, Hui Wang, Yang Xiang, and Yue Yu. 2023. Model Provenance via Model DNA. arXiv:cs.LG/2308.02121
- Mueller et al. (2024) Aaron Mueller, Jannik Brinkmann, Millicent L. Li, Samuel Marks, Koyena Pal, Nikhil Prakash, Can Rager, Aruna Sankaranarayanan, Arnab Sen Sharma, Jiuding Sun, Eric Todd, David Bau, and Yonatan Belinkov. 2024. The Quest for the Right Mediator: A History, Survey, and Theoretical Grounding of Causal Interpretability. CoRR abs/2408.01416 (2024). https://doi.org/10.48550/ARXIV.2408.01416 arXiv:2408.01416
@@
- Navon et al. (2023) Aviv Navon, Aviv Shamsian, Idan Achituve, Ethan Fetaya, Gal Chechik, and Haggai Maron. 2023. Equivariant Architectures for Learning in Deep Weight Spaces. In Proceedings of the 40th International Conference on Machine Learning (Proceedings of Machine Learning Research), Andreas Krause, Emma Brunskill, Kyunghyun Cho, Barbara Engelhardt, Sivan Sabato, and Jonathan Scarlett (Eds.), Vol. 202. PMLR, 25790–25816. https://proceedings.mlr.press/v202/navon23a.html
- OpenAI ([n.d.]) OpenAI. [n.d.]. Chat GPT. https://chat.openai.com/chat.
- OpenAI (2023) OpenAI. 2023. GPT-4 Technical Report. CoRR abs/2303.08774 (2023). https://doi.org/10.48550/ARXIV.2303.08774 arXiv:2303.08774
@@
- Ortega et al. (2018) Pedro A Ortega, Vishal Maini, and DeepMind Safety Team. 2018. Building safe artificial intelligence: specification, robustness, and assurance. DeepMind Safety Research Blog (2018).
- Ouyang et al. (2022) Long Ouyang, Jeff Wu, Xu Jiang, Diogo Almeida, Carroll L. Wainwright, Pamela Mishkin, Chong Zhang, Sandhini Agarwal, Katarina Slama, Alex Ray, John Schulman, Jacob Hilton, Fraser Kelton, Luke Miller, Maddie Simens, Amanda Askell, Peter Welinder, Paul Christiano, Jan Leike, and Ryan Lowe. 2022. Training language models to follow instructions with human feedback. arXiv:cs.CL/2203.02155 https://arxiv.org/abs/2203.02155
@@
- Prabhu et al. (2024) Ameya Prabhu, Vishaal Udandarao, Philip Torr, Matthias Bethge, Adel Bibi, and Samuel Albanie. 2024. Lifelong Benchmarks: Efficient Model Evaluation in an Era of Rapid Progress. arXiv:cs.LG/2402.19472 https://arxiv.org/abs/2402.19472
@@
- Reddi et al. (2019) Vijay Janapa Reddi, Christine Cheng, David Kanter, Peter Mattson, Guenther Schmuelling, Carole-Jean Wu, Brian Anderson, Maximilien Breughe, Mark Charlebois, William Chou, Ramesh Chukka, Cody Coleman, Sam Davis, Pan Deng, Greg Diamos, Jared Duke, Dave Fick, J. Scott Gardner, Itay Hubara, Sachin Idgunji, Thomas B. Jablin, Jeff Jiao, Tom St. John, Pankaj Kanwar, David Lee, Jeffery Liao, Anton Lokhmotov, Francisco Massa, Peng Meng, Paulius Micikevicius, Colin Osborne, Gennady Pekhimenko, Arun Tejusve Raghunath Rajan, Dilip Sequeira, Ashish Sirasao, Fei Sun, Hanlin Tang, Michael Thomson, Frank Wei, Ephrem Wu, Lingjie Xu, Koichi Yamada, Bing Yu, George Yuan, Aaron Zhong, Peizhao Zhang, and Yuchen Zhou. 2019. MLPerf Inference Benchmark. CoRR abs/1911.02549 (2019). arXiv:1911.02549 http://arxiv.org/abs/1911.02549
- Rein et al. (2024) David Rein, Betty Li Hou, Asa Cooper Stickland, Jackson Petty, Richard Yuanzhe Pang, Julien Dirani, Julian Michael, and Samuel R. Bowman. 2024. GPQA: A Graduate-Level Google-Proof Q&A Benchmark. In First Conference on Language Modeling. https://openreview.net/forum?id=Ti67584b98
- Ribeiro et al. (2016) Marco Tulio Ribeiro, Sameer Singh, and Carlos Guestrin. 2016. " Why should i trust you?" Explaining the predictions of any classifier. In Proceedings of the 22nd ACM SIGKDD international conference on knowledge discovery and data mining. 1135–1144.
- Rombach et al. (2022) Robin Rombach, Andreas Blattmann, Dominik Lorenz, Patrick Esser, and Björn Ommer. 2022. High-Resolution Image Synthesis with Latent Diffusion Models. In IEEE/CVF Conference on Computer Vision and Pattern Recognition, CVPR 2022, New Orleans, LA, USA, June 18-24, 2022. IEEE, 10674–10685. https://doi.org/10.1109/CVPR52688.2022.01042
@@
- Ruehle et al. (2021) Victor Ruehle, Robert Sim, Sergey Yekhanin, Nishanth Chandran, Melissa Chase, Daniel Jones, Kim Laine, Boris Köpf, Jaime Teevan, Jim Kleewein, et al. 2021. Privacy preserving machine learning: Maintaining confidentiality and preserving trust. https://www.microsoft.com/en-us/research/blog/privacy-preserving-machine-learning-maintaining-confidentiality-and-preserving-trust/
@@
- Schick et al. (2024) Timo Schick, Jane Dwivedi-Yu, Roberto Dessì, Roberta Raileanu, Maria Lomeli, Eric Hambro, Luke Zettlemoyer, Nicola Cancedda, and Thomas Scialom. 2024. Toolformer: Language models can teach themselves to use tools. Advances in Neural Information Processing Systems 36 (2024).
- Schramowski et al. (2023) Patrick Schramowski, Manuel Brack, Björn Deiseroth, and Kristian Kersting. 2023. Safe latent diffusion: Mitigating inappropriate degeneration in diffusion models. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition. 22522–22531.
- Schürholt et al. (2022) Konstantin Schürholt, Diyar Taskiran, Boris Knyazev, Xavier Giró-i Nieto, and Damian Borth. 2022. Model zoos: A dataset of diverse populations of neural network models. Advances in Neural Information Processing Systems 35 (2022), 38134–38148.
- Schwarzschild et al. (2024) Avi Schwarzschild, Zhili Feng, Pratyush Maini, Zachary C. Lipton, and J. Zico Kolter. 2024. Rethinking LLM Memorization through the Lens of Adversarial Compression. arXiv:cs.LG/2404.15146 https://arxiv.org/abs/2404.15146
@@
- Sharkey et al. (2025) Lee Sharkey, Bilal Chughtai, Joshua Batson, Jack Lindsey, Jeff Wu, Lucius Bushnaq, Nicholas Goldowsky-Dill, Stefan Heimersheim, Alejandro Ortega, Joseph Bloom, Stella Biderman, Adria Garriga-Alonso, Arthur Conmy, Neel Nanda, Jessica Rumbelow, Martin Wattenberg, Nandi Schoots, Joseph Miller, Eric J. Michaud, Stephen Casper, Max Tegmark, William Saunders, David Bau, Eric Todd, Atticus Geiger, Mor Geva, Jesse Hoogland, Daniel Murfet, and Tom McGrath. 2025. Open Problems in Mechanistic Interpretability. arXiv:cs.LG/2501.16496 https://arxiv.org/abs/2501.16496
@@
- Shi et al. (2024) Weijia Shi, Anirudh Ajith, Mengzhou Xia, Yangsibo Huang, Daogao Liu, Terra Blevins, Danqi Chen, and Luke Zettlemoyer. 2024. Detecting Pretraining Data from Large Language Models. In The Twelfth International Conference on Learning Representations. https://openreview.net/forum?id=zWqr3MQuNs
- Shokri et al. (2017) Reza Shokri, Marco Stronati, Congzheng Song, and Vitaly Shmatikov. 2017. Membership Inference Attacks Against Machine Learning Models. In 2017 IEEE Symposium on Security and Privacy (SP). 3–18. https://doi.org/10.1109/SP.2017.41
- Shraga and Miller (2023) Roee Shraga and Renée J. Miller. 2023. Explaining Dataset Changes for Semantic Data Versioning with Explain-Da-V. Proc. VLDB Endow. 16, 6 (2023), 1587–1600. https://doi.org/10.14778/3583140.3583169
@@
- Singh et al. (2024) Chandan Singh, Jeevana Priya Inala, Michel Galley, Rich Caruana, and Jianfeng Gao. 2024. Rethinking Interpretability in the Era of Large Language Models. arXiv:cs.CL/2402.01761
@@
- Spadari et al. (2024) Vincenzo Spadari, Francesco Cerasuolo, Giampaolo Bovenzi, and Antonio Pescapè. 2024. An MLOps Framework for Explainable Network Intrusion Detection with MLflow. In 2024 IEEE Symposium on Computers and Communications (ISCC). 1–6. https://doi.org/10.1109/ISCC61673.2024.10733700
- Staab et al. (2025) Robin Staab, Mark Vero, Mislav Balunovic, and Martin Vechev. 2025. Language Models are Advanced Anonymizers. In The Thirteenth International Conference on Learning Representations. https://openreview.net/forum?id=82p8VHRsaK
@@
- Sun et al. (2025) Mingjie Sun, Yida Yin, Zhiqiu Xu, J. Zico Kolter, and Zhuang Liu. 2025. Idiosyncrasies in Large Language Models. arXiv:cs.CL/2502.12150 https://arxiv.org/abs/2502.12150
@@
- Tabassi (2023) Elham Tabassi. 2023. Artificial Intelligence Risk Management Framework (AI RMF 1.0). https://doi.org/10.6028/NIST.AI.100-1
- Tabassi (2024) Elham Tabassi. 2024. Regulation (EU) 2024/1689 of the European Parliament and of the Council of 13 June 2024 laying down harmonised rules on artificial intelligence and amending Regulations (EC) No 300/2008, (EU) No 167/2013, (EU) No 168/2013, (EU) 2018/858, (EU) 2018/1139 and (EU) 2019/2144 and Directives 2014/90/EU, (EU) 2016/797 and (EU) 2020/1828 (Artificial Intelligence Act) (Text with EEA relevance) PE/24/2024/REV/1, OJ L, 2024/1689, 12.7.2024. https://eur-lex.europa.eu/eli/reg/2024/1689/oj
@@
- Vishwambari and Agrawal (2023) T Vishwambari and Sonali Agrawal. 2023. Integration of Open-Source Machine Learning Operations Tools into a Single Framework. In 2023 International Conference on Computing, Communication, and Intelligent Systems (ICCCIS). 335–340. https://doi.org/10.1109/ICCCIS60361.2023.10425558
- Wang et al. (2024a) Keyu Wang, Abdullah Norozi Iranzad, Scott Schaffter, Doina Precup, and Jonathan Lebensold. 2024a. Mitigating Downstream Model Risks via Model Provenance. arXiv:cs.LG/2410.02230 https://arxiv.org/abs/2410.02230
@@
- Wang et al. (2023) Sheng-Yu Wang, Alexei A Efros, Jun-Yan Zhu, and Richard Zhang. 2023. Evaluating data attribution for text-to-image models. In Proceedings of the IEEE/CVF International Conference on Computer Vision. 7192–7203.
- Wang et al. (2024b) Wenxiao Wang, Weiming Zhuang, and Lingjuan Lyu. 2024b. Towards Fundamentally Scalable Model Selection: Asymptotically Fast Update and Selection. CoRR abs/2406.07536 (2024). https://doi.org/10.48550/ARXIV.2406.07536 arXiv:2406.07536
- Wei et al. (2022) Jason Wei, Yi Tay, Rishi Bommasani, Colin Raffel, Barret Zoph, Sebastian Borgeaud, Dani Yogatama, Maarten Bosma, Denny Zhou, Donald Metzler, Ed H. Chi, Tatsunori Hashimoto, Oriol Vinyals, Percy Liang, Jeff Dean, and William Fedus. 2022. Emergent Abilities of Large Language Models. Transactions on Machine Learning Research (2022). https://openreview.net/forum?id=yzkSU5zdwD Survey Certification.
- Winata et al. (2024) Genta Indra Winata, Hanyang Zhao, Anirban Das, Wenpin Tang, David D. Yao, Shi-Xiong Zhang, and Sambit Sahu. 2024. Preference Tuning with Human Feedback on Language, Speech, and Vision Tasks: A Survey. arXiv:cs.CL/2409.11564 https://arxiv.org/abs/2409.11564
- Xin et al. (2024) Rui Xin, Niloofar Mireshghallah, Shuyue Stella Li, Michael Duan, Hyunwoo Kim, Yejin Choi, Yulia Tsvetkov, Sewoong Oh, and Pang Wei Koh. 2024. A False Sense of Privacy: Evaluating Textual Data Sanitization Beyond Surface-level Privacy Leakage. In Neurips Safe Generative AI Workshop 2024. https://openreview.net/forum?id=3JLtuCozOU
- Yan et al. (2024) Biwei Yan, Kun Li, Minghui Xu, Yueyan Dong, Yue Zhang, Zhaochun Ren, and Xiuzhen Cheng. 2024. On protecting the data privacy of large language models (llms): A survey. arXiv preprint arXiv:2403.05156 (2024).
- Youssef et al. (2024) Paul Youssef, Zhixue Zhao, Jörg Schlötterer, and Christin Seifert. 2024. Detecting Edited Knowledge in Language Models. arXiv:cs.CL/2405.02765 https://arxiv.org/abs/2405.02765
@@
- Zhao et al. (2021) Yan Zhao, Franck Ravat, Julien Aligon, Chantal Soule-dupuy, Gabriel Ferrettini, and Imen Megdiche. 2021. Analysis-oriented Metadata for Data Lakes. In Proceedings of the 25th International Database Engineering & Applications Symposium (IDEAS ’21). Association for Computing Machinery, New York, NY, USA, 194–203. https://doi.org/10.1145/3472163.3472273
- Zhou et al. (2024) Andy Zhou, Bo Li, and Haohan Wang. 2024. Robust Prompt Optimization for Defending Language Models Against Jailbreaking Attacks. In The Thirty-eighth Annual Conference on Neural Information Processing Systems. https://openreview.net/forum?id=jXs6Cvpe7k
- Zhou et al. (2025) Zhanpeng Zhou, Zijun Chen, Yilan Chen, Bo Zhang, and Junchi Yan. 2025. On the emergence of cross-task linearity in pretraining-finetuning paradigm. In Proceedings of the 41st International Conference on Machine Learning (ICML’24). JMLR.org, Article 2559, 31 pages.
@@
- Zhu et al. (2025) Sally Zhu, Ahmed Ahmed, Rohith Kuditipudi, and Percy Liang. 2025. Independence Tests for Language Models. arXiv:cs.LG/2502.12292 https://arxiv.org/abs/2502.12292
- Zou et al. (2023a) Andy Zou, Long Phan, Sarah Chen, James Campbell, Phillip Guo, Richard Ren, Alexander Pan, Xuwang Yin, Mantas Mazeika, Ann-Kathrin Dombrowski, et al. 2023a. Representation engineering: A top-down approach to ai transparency. arXiv preprint arXiv:2310.01405 (2023).
- Zou et al. (2023b) Andy Zou, Zifan Wang, Nicholas Carlini, Milad Nasr, J Zico Kolter, and Matt Fredrikson. 2023b. Universal and transferable adversarial attacks on aligned language models. arXiv preprint arXiv:2307.15043 (2023).